# Exact Fisher zeros and thermofield dynamics across a quantum critical point

### v3

@@ Abstract
By setting the inverse temperature $\beta$ loose to occupy the complex plane, Michael E. Fisher showed that the zeros of the complex partition function $Z$, if approaching the real $\beta$ axis, reveal a thermodynamic phase transition. More recently, Fisher zeros were used to mark the dynamical phase transition in quench dynamics. It remains unclear, however, how Fisher zeros can be employed to better understand quantum phase transitions or the non-unitary dynamics of open quantum systems. Here we answer this question by a comprehensive analysis of the analytically continued one-dimensional transverse field Ising model. We exhaust all the Fisher zeros to show that in the thermodynamic limit they congregate into a remarkably simple pattern in the form of continuous open or closed lines. These Fisher lines evolve smoothly as the coupling constant is tuned, and a qualitative change identifies the quantum critical point. By exploiting the connection between $Z$ and the thermofield double states, we obtain analytical expressions for the short- and long-time dynamics of the survival amplitude including its scaling behavior at the quantum critical point. We point out $Z$ can be realized and probed in monitored quantum circuits. The exact analytical results are corroborated by numerical tensor renormalization group. We further show similar patterns of Fisher zeros also emerge in other spin models. Therefore the approach outlined may serve as a powerful tool for interacting quantum systems.
@@
There is a renewed interest in the concept of complex-valued partition function and its zeros in quantum many-body physics. In the seminal work of Lee and Yang on phase transitions, the magnetic field in spin Hamiltonians (or the chemical potential in a grand canonical ensemble) is analytically continued to take on complex values [1, 2]. Then phase transitions in the thermodynamic limit can be inferred and analyzed by tracking the zeros of the complex-valued partition function $Z$ evaluated for finite-size systems. Fisher extended the recipe by allowing the inverse temperature $\beta=1/k_{B}T$ to be complex instead, which applies to all systems [3]. Yet, compared to the Lee-Yang zeros, the Fisher zeros are more challenging to find or analyze. Because $\beta$ couples to every term in the Hamiltonian, the locations of Fisher zeros on the complex $\beta$ plane appear to feature less regularity. They may not form simple smooth curves as the thermodynamic limit is approached [4], in sharp contrast to the celebrated circle theorem for Lee-Yang zeros [1, 2]. To unveil the unifying patterns behind the ostensible complexity of Fisher zeros, exact analytical results would be greatly helpful.
@@
Figure 1: The continuous lines of Fisher zeros on the complex $\beta=\beta_{r}+i\beta_{i}$ plane for 1DTFIM (schematic). The arrows indicate the velocities of the zeros as the transverse field $g$ is increased. At $g=0$, the zeros congregate into a set of open lines. As $g$ increases, closed lines (ovals) of zeros appear, while the open lines shift upwards to larger $\beta_{r}$ and eventually vanish at the QCP ($g=g_{c}=1$). The closed lines continue to move towards the origin as $g$ is further increased beyond $g_{c}$.
@@
This work is inspired by the confluence of these ideas. By exhausting the Fisher zeros for a canonical quantum many-body system, the 1DTFIM, we describe how insights about quantum phase transitions and quantum dynamics can be gleaned from the complex-valued partition functions ${Z}(\beta)$. A few surprising results are obtained. (1) In the thermodynamic limit, the Fisher zeros of 1DTFIM form continuous curves (lines and loops) which move smoothly as the coupling constant $g$ is tuned. (2) Even though the Fisher zeros never touch the real $\beta$ axis, the quantum critical point can be unambiguously inferred from either the vanishing of open lines (Fig. 1) or the scaling of low-energy excitations extracted from ${Z}$. (3) We exploit the connection between ${Z}$ and the maximally entangled thermofield double (TFD) states [14], which are dual to eternal black holes according to the AdS/CFT correspondence [15], to make analytical predictions about the dynamics of TFD. These include the short-time decay of the survival probability, its periodic oscillation at low temperatures, and the recurrence time at the quantum critical point. (4) We further show that aside from being a useful theoretical device, ${Z}$ can be realized and probed in nonunitary quantum circuits where the unitary evolution of the qubits is interspersed with projective measurements and post selection.
@@
where $\sigma^{x}_{i}$ and $\sigma^{z}_{i}$ are the Pauli matrix operators on site $i$, and the coupling constant $g$ is the transverse magnetic field measured in units of the nearest neighbor Ising interaction (set to 1). As well known, in the thermodynamic limit ($L\rightarrow\infty$), a quantum phase transition at the critical coupling $g_{c}=1$ separates the ferromagnetic phase ($g<1$) from the paramagnetic phase ($g>1$).
@@
Tensor renormalization group with complex $\beta$. We also compute $Z$ using an independent numerical method which serves two purposes. First, it confirms the analytical solution Eq. (2) is correct. Second, it demonstrates a general and precise technique to compute $Z(\beta\in\mathbb{C})$ for interacting quantum spin models for which there is no exact solution. Since $Z$ can be represented as the trace of a product of local tensors, it can be evaluated by tensor network algorithms [23, 24, 25] which have proven to be efficient and accurate for spin models [26]. Starting from the 1D system, we first construct a 2D “lattice” of size $L\times N$ by Trotter decomposition $\beta=\tau N$. Then, $Z$ is obtained from High Order Tensor Renormalization Group (HOTRG) [27]. We have checked that tensor results are in agreement with both exact diagonalization and the exact solution above (see [22] for details).
@@
It is worth to mention that when $\beta$ is complex, Monte Carlo develops sign problems but HOTRG can yield $Z$ accurately [26]. Previously renormalization group (RG) transformation has been generalized to the complex $\beta$ plane to study confinement in lattice gauge theory and classical $O(N)$ models. It was observed that the Fisher zeros are located at the boundary of the attraction basins of infrared fixed points, and they control the global properties of the complex RG flows [28, 29]. Here we confine our attention to the evaluation of Fisher zeros through HOTRG but for quantum spin models.
@@
Motion of Fisher zeros across QCP. The exact solution corroborated by HOTRG yields a complete picture, sketched in Fig. 1, of the evolution of Fisher zeros on the complex $\beta$ plane as the coupling constant $g$ is varied. The main observations are summarized into (a) to (d) below. (a) The Fisher zeros of 1DTFIM appear as continuous lines, rather than isolated points or densely packed regions, in the thermodynamic limit $L\rightarrow\infty$. (b) For $g<1$, there exist open lines of zeros that are approximately parallel to the $\beta_{r}$ axis. (c) As $g$ increases from 0 to 1, these open lines shift upwards toward large $\beta_{i}$, and vanish completely at the QCP $g=1$. (d) Away from $g=0$, closed loops of Fisher zeros also appear and persist for all $g$. These loops are centered around the $\beta_{i}$ axis and move towards $\beta_{i}=0$ as $g$ is increased. We stress that properties (a) to (d) came as surprises, in the sense that they are not obvious from staring at the model or the exact solution of $Z$.
@@
Some of these features can be appreciated by considering a few limit cases. At $g=0$, $Z=\mathrm{Tr}V^{L}$ where the transfer matrix $V=e^{\beta}+\sigma_{x}e^{-\beta}$. The Fisher zeros are determined by the condition $\tanh\beta=e^{in\pi/L}$ with $n$ an integer, i.e. they are uniformly distributed on the unit circle on the plane of $\tanh\beta$ [13]. On the complex $\beta$ plane, the zeros form a set of straight lines at $\beta_{i}=\pi/4+m\pi/2$, with $m$ an integer, as $L\rightarrow\infty$. Another simple case is the small size system with $L=2$. After some algebra, one finds the zeros are located on the $\beta_{i}$ axis at
@@
It is clear from (c) that the QCP is marked by a qualitative change in the Fisher zeros. To describe the lowest open Fisher line, one can take its imaginary part $\beta_{i}$ at some fixed large value of $\beta_{r}$ and plot $1/\beta_{i}$ against $g$. In Ref. [5], it was shown that the resulting curve is linear near QCP. Similarly, for the closed Fisher lines, we can select a characteristic point on each curve, e.g. recording the $\beta_{i}$ value of the rightmost point. Then a linear relationship between $1/\beta_{i}$ and $g$ is observed on the quantum disordered side $g>1$, which is consistent with the spin flip excitations. A more refined description is to define the velocities of the zeros on the complex $\beta$ plane with increasing $g$. Interestingly, we find the speed of the zeros exhibits properties similar to the Grüneisen parameter [31, 32], see [22].
@@
Fisher zeros in other spin models. We conjecture that the regularity of Fisher’s zeros and its correlation with QCP, observed above for 1DTFIM, generalize to a host of other systems. Since ${Z}$ and its zeros can be reliably computed from tensor network method, the program outlined here can serve as a powerful diagnostic tool for quantum many-body physics. To illustrate this, we consider two examples. The first is the axial next nearest neighbour Ising (ANNNI) model with the Hamiltonian
@@
where we chose $\gamma=0.01$, bringing the system close to a QCP. For both systems, numerical calculations show that as $L$ increases, the zeros indeed approach continuous lines. Figure 2 compares the Fisher zeros of these two models obtained from HOTRG with $L=32$ [22]. For the ANNNI model, open lines corresponding to domain wall excitations still exist. But the zero locations are no longer symmetric between $\beta_{r}>0$ and $\beta_{r}<0$, causing the open lines to tilt. For the XY model near QCP, we witness the same qualitative change as seen in the 1DTFIM: the open lines completely vanish, giving way to closed loops. The change is mandatory because otherwise, the two phases separated by the QCP would share the same global patterns of RG flows on the complex $\beta$ plane, which would contradict to the fact that they correspond to different RG fixed points.
What can $Z$ tell us about quantum dynamics? We now view $Z$ from another perspective, as the survival amplitude of certain quantum state. Let $\beta=\beta_{r}+it$, and interpret the imaginary (real) part of $\beta$ as time $t$ (the inverse temperature $\beta_{r}=1/k_{B}T$). Let $E_{n}$ be the eigen energies, and $|n\rangle$ the corresponding eigen states, of Hamiltonian $H$. Then
@@
$S(\beta_{r},t)=\left|\frac{Z(\beta_{r},t)}{Z(\beta_{r},0)}\right|^{2}$
which is normalized to 1 as $t=0$. The disorder average of $S$ has been used to diagnose chaos and information scrambling in Hamiltonians containing random couplings, e.g. the SYK model [33]. Following Ref. [34], we purify the mixed state above by introducing the thermofield double state in an enlarged Hilbert space
$|\psi(\beta_{r},0)\rangle=\frac{1}{\sqrt{Z(\beta_{r},0)}}\sum_{n}e^{-\beta_{r}E_{n}/2}|n\rangle_{L}\otimes|n\rangle_{R}$
where for clarity the subscripts $L,R$ are used to denote the left and right copy respectively [14]. Under the single Hamiltonian $H_{L}\otimes 1_{R}$, the TFD state evolves with time as $\psi(\beta_{r},t)=e^{-it(H_{L}\otimes 1_{R})}|\psi(\beta_{r},0)\rangle$. One finds that $S$ is nothing but the survival probability [34]
$S(\beta_{r},t)=|\langle\psi(\beta_{r},t)|\psi(\beta_{r},0)\rangle|^{2}$
@@
The short time dynamics of $S$ is characterized by exponential decay $S(t)\simeq e^{-(t/\tau)^{2}}$ with $\tau^{2}={k_{B}\beta_{r}/C_{V}}$. Fig. 3(a) shows the specific heat capacity $C_{V}/L$ obtained by fitting $S(t)$ at some typical values of $g$. The decay rate $1/\tau\propto\sqrt{C_{V}}$ correlates with the locations of Fisher zeros at which $S$ is forced to vanish. This can be seen on the low $\beta_{r}$ portion of Fig. 3(a): as $g$ is increased from $g=1$, the Fisher zeros move closer to the $\beta_{r}$ axis, and accordingly $S$ is suppressed from $1$ to zero at an increasing rate. At high $\beta_{r}$ (low temperatures), however, the decay is most rapid at the QCP, where quantum fluctuation is strong and $C_{V}$ reaches maximum.
@@
$S(t)\simeq 1+2e^{-\beta_{r}\Delta}\cos(\Delta t)+...$
@@
The dynamics of $S$ exhibits rich scaling behaviors at the QCP. For a system of finite size $L$, the energy spectrum is discrete and the initial TFD state will come back to itself after the quantum recurrence time $t^{*}$. To find $t^{*}$, take the limit of large $\beta_{r}$, expand the exact Suzuki solution and carry out the $k$-sum, we find at $g=1$ the oscillation frequency is $\Delta=2\tan({\pi}/{4L})$ which becomes ${\pi}/{2L}$ in the limit of large $L$ to yield $t^{*}=4L$. For smaller $\beta_{r}$, $S(t)$ deviates significantly from the sinusoidal form to acquire fine structures, but $t^{*}$ stays approximately the same [22]. Our prediction $t^{*}=4L$ agrees with the conformal field theory result from Ref [16] (see [22] for details). We also observe that $S$ exhibits self-similarity, i.e. $S(\beta_{r},t)_{L}\approx S(n\beta_{r},nt)_{nL}$ with $n$ an integer, at the QCP and for intermediate to large $(\beta_{r},t)$. Figure. 3(c) shows the example of $n=2$. This is consistent with the notion of criticality and provides a nice way to define RG flow in the complex plane (see [22] for details).
Probing $Z$ from monitored quantum circuits. Lastly we show that $Z$ is not just a formal theoretical construct, it can be realized and probed in quantum circuits. Consider an array of qubits, $j=1,2,...,L$. The standard single- and two-qubit Pauli gates implement unitary evolution such as $U_{zz}(t)=\exp({it\sum_{j}\sigma^{z}_{j}\sigma^{z}_{j+1}})$ and $U_{x}(t)=\exp({itg\sum_{j}\sigma^{x}_{j}})$. Let us define the cycle operator $U_{F}(t)=U_{zz}(t)U_{x}(t)$, then repeated application of $U_{F}$ produces the time evolution corresponding to the 1DTFIM Hamiltonian, albeit in discretized form $U(t)=[U_{F}(t/N)]^{N}$. Now let us introduce nonunitary gates in the form of $U_{x}(i\beta_{r})$ and $U_{zz}(i\beta_{r})$, i.e. by replacing $t$ with the imaginary time $i\beta_{r}$. These gates can be realized by including ancilla qubits and performing projective measurements and post-selection on the ancilla [35, 36, 37], where the value of $\beta_{r}$ characterizes the strength of the measurement. If we intersperse these unitary and nonunitary gates [8], e.g. by forming a cycle
@@
The circuit construction here differs from the approach in Ref. [13] based on the transfer matrix. From an alternative perspective, the nonunitary gates introduce imaginary parts to the coupling constants in the Hamiltonian and render it non-hermitian [38]. Previous studies on non-hermitian Ising-type models mostly focused on their entanglement or spectral properties [35, 36, 37]. Given the recent success in observing Lee-Yang zeros [39] and Loschmidt ratios [40], it seems promising that the circuit connection to $Z$ and TFD states [41] can led to future experimental studies of Fisher zeros [42].
In summary, this work presents a comprehensive analysis of the Fisher zeros in the 1DTFIM. We have demonstrated that Fisher zeros form continuous, smoothly shifting curves in the complex $\beta$ plane, and they offer insights into quantum criticality despite never touching the real $\beta$ axis. The complex partition function $Z$ provides valuable analytical predictions about the quantum dynamics which is beyond the reach of the traditional thermodynamics, and can be realized and probed in nonunitary quantum circuits. Our improvement/clarification over the original Suzuki solution correctly captures the different excitations on the two sides of the QPT and the breakdown of the Kramers-Wannier duality. Successful generalization of the numerical analysis to other (the ANNNI and XY) spin models demonstrates that this approach can open up a promising avenue to study quantum critical systems in 1D [43, 44, 45] and higher dimensions [46, 47, 48].
We thank Youjin Deng, Shijie Hu, Ian McCulloch, Yannick Meurice, Tao Xiang, and Jiahao Yang for helpful discussions. This work is supported by National Natural Science Foundation of China (Grant No. 12274126). EZ acknowledges the support from NSF grant PHY-206419 and AFOSR grant FA9550-23-1-0598.

### v4

@@ Abstract
By setting the inverse temperature $\beta$ loose to occupy the complex plane, Fisher showed that the zeros of the complex partition function $Z$, if approaching the real $\beta$ axis, reveal a thermodynamic phase transition. More recently, Fisher zeros were used to mark the dynamical phase transition in quench dynamics. It remains unclear, however, how Fisher zeros can be employed to better understand quantum phase transitions or the non-unitary dynamics of open quantum systems. Here we answer this question by a comprehensive analysis of the analytically continued one-dimensional transverse field Ising model. We exhaust all the Fisher zeros to show that in the thermodynamic limit they congregate into a remarkably simple pattern in the form of continuous open or closed lines. These Fisher lines evolve smoothly as the coupling constant is tuned, and a qualitative change identifies the quantum critical point. By exploiting the connection between $Z$ and the thermofield double states, we obtain analytical expressions for the short- and long-time dynamics of the survival amplitude, including its scaling behavior at the quantum critical point. We point out $Z$ can be realized and probed in monitored quantum circuits. The exact analytical results are corroborated by the numerical tensor renormalization group. We further show that similar patterns of Fisher zeros also emerge in other spin models. Therefore, the approach outlined may serve as a powerful tool for interacting quantum systems.
@@
Introduction. There is a renewed interest in the concept of complex-valued partition function and its zeros in quantum many-body physics. In the seminal work of Lee and Yang on phase transitions, the magnetic field in spin Hamiltonians (or the chemical potential in a grand canonical ensemble) is analytically continued to take on complex values [1, 2]. Then phase transitions in the thermodynamic limit can be inferred and analyzed by tracking the zeros of the complex-valued partition function $Z$ evaluated for finite-size systems. Fisher extended the recipe by allowing the inverse temperature $\beta=1/k_{B}T$ to be complex instead, which applies to all systems [3]. Yet, compared to the Lee-Yang zeros, the Fisher zeros are more challenging to find or analyze. Because $\beta$ couples to every term in the Hamiltonian, the locations of Fisher zeros on the complex $\beta$ plane appear to feature less regularity. They may not form simple smooth curves as the thermodynamic limit is approached [4], in sharp contrast to the celebrated circle theorem for Lee-Yang zeros [1, 2, 5]. To unveil the unifying patterns behind the ostensible complexity of Fisher zeros, exact analytical results would be greatly helpful.
@@
Figure 1: The continuous lines of Fisher zeros on the complex $\beta=\beta_{r}+i\beta_{i}$ plane for 1DTFIM (schematic). The arrows indicate the velocities of the zeros as the transverse field $g$ is increased. At $g=0$, the zeros congregate into a set of open lines. As $g$ increases, closed lines (ovals) of zeros appear, while the open lines shift upwards to larger $\beta_{r}$ and eventually vanish at the QCP ($g=g_{c}=1$). The closed lines continue to move toward the origin as $g$ is further increased beyond $g_{c}$.
@@
This work is inspired by the confluence of these ideas. By exhausting the Fisher zeros for a canonical quantum many-body system, the 1DTFIM, we describe how insights about quantum phase transitions and quantum dynamics can be gleaned from the complex-valued partition functions ${Z}(\beta)$. A few surprising results are obtained. (i) In the thermodynamic limit, the Fisher zeros of 1DTFIM form continuous curves (lines and loops) which move smoothly as the coupling constant $g$ is tuned. (ii) Even though the Fisher zeros never touch the real $\beta$ axis, the quantum critical point can be unambiguously inferred from either the vanishing of open lines (Fig. 1) or the scaling of low-energy excitations extracted from ${Z}$. (iii) We exploit the connection between ${Z}$ and the maximally entangled thermofield double (TFD) states [15], which are dual to eternal black holes according to the AdS/CFT correspondence [16], to make analytical predictions about the dynamics of TFD. These include the short-time decay of the survival probability, its periodic oscillation at low temperatures, and the recurrence time at the quantum critical point. (iv) We further show that aside from being a useful theoretical device, ${Z}$ can be realized and probed in nonunitary quantum circuits where the unitary evolution of the qubits is interspersed with projective measurements and post selection.
@@
where $\sigma^{x}_{i}$ and $\sigma^{z}_{i}$ are the Pauli matrix operators on site $i$, and the coupling constant $g$ is the transverse magnetic field measured in units of the nearest neighbor Ising interaction (set to one). As is well known, in the thermodynamic limit ($L\rightarrow\infty$), a quantum phase transition at the critical coupling $g_{c}=1$ separates the ferromagnetic phase ($g<1$) from the paramagnetic phase ($g>1$).
@@
Tensor renormalization group with complex $\beta$. We also compute $Z$ using an independent numerical method which serves two purposes. First, it confirms the analytical solution Eq. (2) is correct. Second, it demonstrates a general and precise technique to compute $Z(\beta\in\mathbb{C})$ for interacting quantum spin models for which there is no exact solution. Since $Z$ can be represented as the trace of a product of local tensors, it can be evaluated by tensor network algorithms [24, 25, 26] which have proven to be efficient and accurate for spin models [27]. Starting from the 1D system, we first construct a 2D “lattice” of size $L\times N$ by Trotter decomposition $\beta=\tau N$. Then, $Z$ is obtained from the high order tensor renormalization group (HOTRG) [28]. We have checked that tensor results are in agreement with both exact diagonalization and the exact solution above (see [23] for details).
@@
It is worth to mention that when $\beta$ is complex, Monte Carlo develops sign problems but HOTRG can yield $Z$ accurately [27]. Previously, renormalization group (RG) transformation has been generalized to the complex $\beta$ plane to study confinement in lattice gauge theory and classical $O(N)$ models. It was observed that the Fisher zeros are located at the boundary of the attraction basins of infrared fixed points, and they control the global properties of the complex RG flows [29, 30]. Here we confine our attention to the evaluation of Fisher zeros through HOTRG but for quantum spin models.
@@
Motion of Fisher zeros across QCP. The exact solution corroborated by HOTRG yields a complete picture, sketched in Fig. 1, of the evolution of Fisher zeros on the complex $\beta$ plane as the coupling constant $g$ is varied. The main observations are summarized into (i) to (iv) below. (i) The Fisher zeros of 1DTFIM appear as continuous lines, rather than isolated points or densely packed regions, in the thermodynamic limit $L\rightarrow\infty$. (ii) For $g<1$, there exist open lines of zeros that are approximately parallel to the $\beta_{r}$ axis. (iii) As $g$ increases from zero to one, these open lines shift upwards toward large $\beta_{i}$, and vanish completely at the QCP $g=1$. (iv) Away from $g=0$, closed loops of Fisher zeros also appear and persist for all $g$. These loops are centered around the $\beta_{i}$ axis and move toward $\beta_{i}=0$ as $g$ is increased. We stress that properties (i) to (iv) came as surprises, in the sense that they are not obvious from staring at the model or the exact solution of $Z$.
@@
Some of these features can be appreciated by considering a few limit cases. At $g=0$, $Z=\mathrm{Tr}V^{L}$ where the transfer matrix $V=e^{\beta}+\sigma_{x}e^{-\beta}$. The Fisher zeros are determined by the condition $\tanh\beta=e^{in\pi/L}$ with $n$ an integer, i.e., they are uniformly distributed on the unit circle on the plane of $\tanh\beta$ [14]. On the complex $\beta$ plane, the zeros form a set of straight lines at $\beta_{i}=\pi/4+m\pi/2$, with $m$ an integer, as $L\rightarrow\infty$. Another simple case is the small size system with $L=2$. After some algebra, one finds the zeros are located on the $\beta_{i}$ axis at
@@
It is clear from (iii) that the QCP is marked by a qualitative change in the Fisher zeros. To describe the lowest open Fisher line, one can take its imaginary part $\beta_{i}$ at some fixed large value of $\beta_{r}$ and plot $1/\beta_{i}$ against $g$. In Ref. [6], it was shown that the resulting curve is linear near QCP. Similarly, for the closed Fisher lines, we can select a characteristic point on each curve, e.g., recording the $\beta_{i}$ value of the rightmost point. Then a linear relationship between $1/\beta_{i}$ and $g$ is observed on the quantum disordered side $g>1$, which is consistent with the spin flip excitations. A more refined description is to define the velocities of the zeros on the complex $\beta$ plane with increasing $g$. Interestingly, we find the speed of the zeros exhibits properties similar to the Grüneisen parameter [32, 33], see [23].
@@
Fisher zeros in other spin models. We conjecture that the regularity of Fisher’s zeros and its correlation with QCP, observed above for 1DTFIM, generalize to a host of other systems. Since ${Z}$ and its zeros can be reliably computed from the tensor network method, the program outlined here can serve as a powerful diagnostic tool for quantum many-body physics. To illustrate this, we consider two examples. The first is the axial next nearest neighbor Ising (ANNNI) model with the Hamiltonian
@@
where we chose $\gamma=0.01$, bringing the system close to a QCP. For both systems, numerical calculations show that as $L$ increases, the zeros indeed approach continuous lines. Figure 2 compares the Fisher zeros of these two models obtained from HOTRG with $L=32$ [23]. For the ANNNI model, open lines corresponding to domain wall excitations still exist. But the zero locations are no longer symmetric between $\beta_{r}>0$ and $\beta_{r}<0$, causing the open lines to tilt. For the XY model near QCP, we witness a similar qualitative change as seen in the 1DTFIM: the open lines completely vanish, giving way to closed loops. The change is mandatory because otherwise, the two phases separated by the QCP would share the same global patterns of RG flows on the complex $\beta$ plane, which would contradict the fact that they correspond to different RG fixed points.
What can $Z$ tell us about quantum dynamics? We now view $Z$ from another perspective, as the survival amplitude of a certain quantum state. Let $\beta=\beta_{r}+it$, and interpret the imaginary (real) part of $\beta$ as time $t$ (the inverse temperature $\beta_{r}=1/k_{B}T$). Let $E_{n}$ be the eigen energies, and $|n\rangle$ the corresponding eigen states, of Hamiltonian $H$. Then
@@
$S(\beta_{r},t)=\left|\frac{Z(\beta_{r},t)}{Z(\beta_{r},0)}\right|^{2},$
which is normalized to 1 as $t=0$. The disorder average of $S$ has been used to diagnose chaos and information scrambling in Hamiltonians containing random couplings, e.g., the SYK model [34]. Following Ref. [35], we purify the mixed state above by introducing the thermofield double state in an enlarged Hilbert space
$|\psi(\beta_{r},0)\rangle=\frac{1}{\sqrt{Z(\beta_{r},0)}}\sum_{n}e^{-\beta_{r}E_{n}/2}|n\rangle_{L}\otimes|n\rangle_{R},$
where, for clarity, the subscripts $L,R$ are used to denote the left and right copy, respectively [15]. Under the single Hamiltonian $H_{L}\otimes 1_{R}$, the TFD state evolves with time as $\psi(\beta_{r},t)=e^{-it(H_{L}\otimes 1_{R})}|\psi(\beta_{r},0)\rangle$. One finds that $S$ is nothing but the survival probability [35]
$S(\beta_{r},t)=|\langle\psi(\beta_{r},t)|\psi(\beta_{r},0)\rangle|^{2}.$
@@
The short time dynamics of $S$ is characterized by exponential decay $S(t)\simeq e^{-(t/\tau)^{2}}$ with $\tau^{2}={k_{B}\beta_{r}/C_{V}}$. Figure 3(a) shows the specific heat capacity $C_{V}/L$ obtained by fitting $S(t)$ at some typical values of $g$. The decay rate $1/\tau\propto\sqrt{C_{V}}$ correlates with the locations of Fisher zeros at which $S$ is forced to vanish. This can be seen on the low $\beta_{r}$ portion of Fig. 3(a): as $g$ is increased from $g=1$, the Fisher zeros move closer to the $\beta_{r}$ axis, and accordingly $S$ is suppressed from one to zero at an increasing rate. At high $\beta_{r}$ (low temperatures), however, the decay is most rapid at the QCP, where quantum fluctuation is strong and $C_{V}$ reaches maximum.
@@
$S(t)\simeq 1+2e^{-\beta_{r}\Delta}\cos(\Delta t)+...,$
@@
The dynamics of $S$ exhibits rich scaling behaviors at the QCP. For a system of finite size $L$, the energy spectrum is discrete and the initial TFD state will come back to itself after the quantum recurrence time $t^{*}$. To find $t^{*}$, take the limit of large $\beta_{r}$, expand the exact Suzuki solution, and carry out the $k$-sum, we find at $g=1$ the oscillation frequency is $\Delta=2\tan({\pi}/{4L})$ which becomes ${\pi}/{2L}$ in the limit of large $L$ to yield $t^{*}=4L$. For smaller $\beta_{r}$, $S(t)$ deviates significantly from the sinusoidal form to acquire fine structures, but $t^{*}$ stays approximately the same [23]. Our prediction $t^{*}=4L$ agrees with the conformal field theory result from Ref. [17] (see [23] for details). We also observe that $S$ exhibits self-similarity, i.e., $S(\beta_{r},t)_{L}\approx S(n\beta_{r},nt)_{nL}$ with $n$ an integer, at the QCP and for intermediate to large $(\beta_{r},t)$. Figure. 3(c) shows the example of $n=2$. This is consistent with the notion of criticality and provides a nice way to define RG flow in the complex plane (see [23] for details).
Probing $Z$ from monitored quantum circuits. Lastly, we show that $Z$ is not just a formal theoretical construct, it can be realized and probed in quantum circuits. Consider an array of qubits, $j=1,2,...,L$. The standard single- and two-qubit Pauli gates implement unitary evolution such as $U_{zz}(t)=\exp({it\sum_{j}\sigma^{z}_{j}\sigma^{z}_{j+1}})$ and $U_{x}(t)=\exp({itg\sum_{j}\sigma^{x}_{j}})$. Let us define the cycle operator $U_{F}(t)=U_{zz}(t)U_{x}(t)$, then repeated application of $U_{F}$ produces the time evolution corresponding to the 1DTFIM Hamiltonian, albeit in discretized form $U(t)=[U_{F}(t/N)]^{N}$. Now let us introduce nonunitary gates in the form of $U_{x}(i\beta_{r})$ and $U_{zz}(i\beta_{r})$, i.e., by replacing $t$ with the imaginary time $i\beta_{r}$. These gates can be realized by including ancilla qubits and performing projective measurements and post-selection on the ancilla [36, 37, 38], where the value of $\beta_{r}$ characterizes the strength of the measurement. If we intersperse these unitary and nonunitary gates [9], e.g., by forming a cycle
@@
The circuit construction here differs from the approach in Ref. [14] based on the transfer matrix. From an alternative perspective, the nonunitary gates introduce imaginary parts to the coupling constants in the Hamiltonian and render it non-Hermitian [39]. Previous studies on non-Hermitian Ising-type models mostly focused on their entanglement or spectral properties [36, 37, 38]. Given the recent success in observing Lee-Yang zeros [40] and Loschmidt ratios [41], it seems promising that the circuit connection to $Z$ and TFD states [42] can lead to future experimental studies of Fisher zeros [43].
Conclusion. In summary, this work presents a comprehensive analysis of the Fisher zeros in the 1DTFIM. We have demonstrated that Fisher zeros form continuous, smoothly shifting curves in the complex $\beta$ plane, and they offer insights into quantum criticality despite never touching the real $\beta$ axis. The complex partition function $Z$ provides valuable analytical predictions about the quantum dynamics which is beyond the reach of the traditional thermodynamics, and can be realized and probed in nonunitary quantum circuits. Our improvement/clarification over the original Suzuki solution correctly captures the different excitations on the two sides of the QPT and the breakdown of the Kramers-Wannier duality. Successful generalization of the numerical analysis to other (the ANNNI and XY) spin models demonstrates that this approach can open up a promising avenue to study quantum critical systems in 1D [44, 45, 46] and higher dimensions [47, 48, 49].
Acknowledgments. We thank Youjin Deng, Shijie Hu, Ian McCulloch, Yannick Meurice, Tao Xiang, and Jiahao Yang for helpful discussions. This work is supported by the National Natural Science Foundation of China (Grant No. 12274126). E.Z. acknowledges the support from NSF Grant PHY-206419, and AFOSR Grant FA9550-23-1-0598.